OpenStudy (anonymous):

Can anyone please explain how to solve these for x? A.(3x + 24) : 3x = 85 : 51 B.(3x + 24) : 81 = 3x : 51 C.(3x + 24) : 51 = 3x : 85 D.34 : 24 = 3x : 51

OpenStudy (ghazi):

\[\frac{ 3x+24 }{ 3x }= \frac{ 85 }{ 51 }\]

OpenStudy (ghazi):

using cross multiplication 51*(3x+24)= 85*(3x)...just solve for x

OpenStudy (ghazi):

can you do it now?

OpenStudy (anonymous):

Awesome! I was doing it like that at first but I thought I was wrong. Thanks! :D

OpenStudy (ghazi):

:)